A researcher develops a regression equation to predict first-year college grades (Y) from high school GPA (X): Y = 1.46 + .51X + e. A student with a high school GPA of 2.8 had an actual first-year college GPA of 3.7. What was the error of prediction for this student?

A.) 0.5

B.) -0.5

C.) -0.8

D.) 0.8
Answer #1

Solution :

Y = 1.46 + 0.51 X

Plug X = 3.7 we get ,

Y = 1.46 + 0.51 * 2.8 = 2.888

Error prediction = Actual value - predicted value

= 3.7 - 2.888

= 0.8

D)